# Quantifying non-Hermiticity using single- and many-particle quantum properties

### v1

@@ Abstract
The non-Hermitian paradigm of quantum systems displays salient features drastically different from Hermitian counterparts. In this work, we focus on one such aspect, the difference of evolving quantum ensembles under $H_{\mathrm{nh}}$ (right ensemble) versus its Hermitian conjugate, $H_{\mathrm{nh}}^{\dagger}$ (left ensemble). We propose a formalism that quantifies the (dis-)similarity of these right and left ensembles, for single- as well as many-particle quantum properties. Such a comparison gives us a scope to measure the extent to which non-Hermiticity gets translated from the Hamiltonian into physically observable properties. We test the formalism in two cases: First, we construct a non-Hermitian Hamiltonian using a set of imperfect Bell states, showing that the non-Hermiticity of the Hamiltonian does not automatically comply with the non-Hermiticity at the level of observables. Second, we study the interacting Hatano–Nelson model with asymmetric hopping as a paradigmatic quantum many-body Hamiltonian. Interestingly, we identify situations where the measures of non-Hermiticity computed for the Hamiltonian, for single-, and for many-particle quantum properties behave distinctly from each other. Thus, different notions of non-Hermiticity can become useful in different physical scenarios. Furthermore, we demonstrate that the measures can mark the model’s Parity–Time (PT) symmetry-breaking transition. Our findings can be instrumental in unveiling new exotic quantum phases of non-Hermitian quantum many-body systems as well as in preparing resourceful states for quantum technologies.
Figure 1: Behavior of degree of non-Hermiticity computed using $\mathcal{D}=\frac{||H_{\mathrm{nh}}-H_{\mathrm{nh}}^{\dagger}||}{||H_{\mathrm{nh}}||}$ for two different non-Hermitian Hamiltonians. (a) Behavior of $\mathcal{D}$ for $H^{\mathrm{Bell}}_{\mathrm{nh}}$ constructed using a set of imperfect Bell states as defined in Eqs. (16) and (17). For each value of the non-Hermiticity parameter $\bar{\alpha}=1-\alpha$, we plot the maximum ($\mathcal{D}_{\max}$, blue) and the minimum ($\mathcal{D}_{\min}$, red) within 1000 random realizations of the Hamiltonian obtained by choosing its energy eigenvalues $\{\lambda_{m}\}$ from a Gaussian distribution with zero mean and unit variance. While $\mathcal{D}_{\min}$ increases only slowly and under strong fluctuations, $\mathcal{D}_{\max}$ exhibits a clear monotonic growth with $\bar{\alpha}$. (b) Behavior of $\mathcal{D}$ for the interacting non-Hermitian Hatano–Nelson model ($H^{\mathrm{HN}}_{\mathrm{nh}}$). At low values of interaction $V$, $\mathcal{D}$ increases monotonically with asymmetric coupling $\chi$ and saturates to $\mathcal{D}\approx 2$. However, even for moderately large values of $V$, $\mathcal{D}$ takes a significant non-zero value only beyond a threshold coupling, $\chi_{c}$. Further, beyond $V\gtrsim 10^{3}$, $\mathcal{D}$ remains zero for all values of $\chi$ considered in our analysis. In other words, under strong interactions the distance between $H_{\mathrm{nh}}^{\mathrm{HN}}$ and $(H_{\mathrm{nh}}^{\mathrm{HN}})^{\dagger}$ becomes insignificant in comparison to the norm of $H_{\mathrm{nh}}^{\mathrm{HN}}$ (see also Fig. 6 in Appendix C). The white circles correspond to the PT symmetry-breaking transition of the model marked by non-analytic behavior of the finite-size level-spacing $\Delta_{01}:=\mathrm{Re}(\lambda_{1}-\lambda_{0})$ (inset: $\Delta_{01}$ versus $V$ for fixed $\chi=2.5$). There is a striking agreement between the regions where $\mathcal{D}$ drops and non-analyticities in $\Delta_{01}$. The data is reported for $N=12$ sites, half-filling, and with anti-periodic boundary conditions.
@@
Here, we present such a framework. To provide a concise overview, we briefly anticipate its main features, while we discuss its details in the bulk of the article. As is well known, the left and right eigenvectors diagonalizing a non-Hermitian Hamiltonian $H_{\mathrm{nh}}=\sum_{m}\lambda_{m}|R_{m}\rangle\langle L_{m}|$, obeying $H_{\mathrm{nh}}|R_{m}\rangle=\lambda_{m}|R_{m}\rangle$, $\langle L_{m}|H_{\mathrm{nh}}=\langle L_{m}|\lambda_{m}$, and $\langle L_{m}|R_{n}\rangle=\delta_{mn}$, can in general be distinct, $\langle L_{m}|\neq(|R_{m}\rangle)^{\dagger}$. As we discuss, the evolution generated by $H_{\mathrm{nh}}$ leads an initial state to ensembles of right eigenvectors (including, e.g., single eigenstates reachable at large times), $\rho_{RR}(t)=\frac{e^{-itH_{\mathrm{nh}}}\rho_{\mathrm{in}}e^{itH^{\dagger}_{\mathrm{nh}}}}{\mathrm{Tr}[e^{-itH_{\mathrm{nh}}}\rho_{\mathrm{in}}e^{itH^{\dagger}_{\mathrm{nh}}}]}$, such that observables measured on the system will physically be of the form $\langle R|\hat{O}|R\rangle$. In contrast, the evolution governed by $H_{\mathrm{nh}}^{\dagger}$ leads to ensembles of left eigenvectors, $\rho_{LL}(t)=\frac{e^{-itH^{\dagger}_{\mathrm{nh}}}\rho_{\mathrm{in}}e^{itH_{\mathrm{nh}}}}{\mathrm{Tr}[e^{-itH^{\dagger}_{\mathrm{nh}}}\rho_{\mathrm{in}}e^{itH_{\mathrm{nh}}}]}$, and thus to expectation values of the form $\langle L|\hat{O}|L\rangle$. As we further show, the dynamics under any non-Hermitian Hamiltonian $H_{\mathrm{nh}}$ can be modeled as an effect of complete quantum measurements and subsequent post-selection of the results, giving a clear physical meaning to these expectation values in terms of observable properties. Unlike the Hermitian case, the evolution governed by $H_{\mathrm{nh}}$ and $H_{\mathrm{nh}}^{\dagger}$ may lead to results that can significantly differ from each other, due to a rich interplay between potentially complex $\lambda_{m}$ and/or the difference between $\{|R_{m}\rangle\}$ and $\{\langle L_{m}|\}$. Our aim is to design measures that systematically capture these differences between $H_{\mathrm{nh}}$ and $H_{\mathrm{nh}}^{\dagger}$, and thus quantify the non-Hermiticity of the system through observable properties.
As a first measure, we introduce the “Hamiltonian non-Hermiticity” $\mathcal{D}=\frac{||H_{\mathrm{nh}}-H_{\mathrm{nh}}^{\dagger}||}{||H_{\mathrm{nh}}||}$, where ||.||||.||| | . | | is the operator norm. This quantifier relies on knowledge of the full model Hamiltonian. In order to make non-Hermiticity more easily accessible, we further define the “non-Hermiticity score” as ${SC}^{\mathcal{F}}_{\mathrm{nh}}=|\mathcal{F}_{RR}[\rho_{RR}]-\mathcal{F}_{LL}[\rho_{LL}]|$. This score quantifies the non-Hermiticity at the level of any physical state-dependent function $\mathcal{F}[.]$ related to the model by measuring the difference when evaluating $\mathcal{F}[.]$ for the state evolved under $H_{\mathrm{nh}}$ versus the one evolved under $H_{\mathrm{nh}}^{\dagger}$.
@@
The above formalism opens up the possibility to analyze to what extent the degree of non-Hermiticity of the Hamiltonian $H_{\mathrm{nh}}$ gets translated to the behavior of physically observable quantities. Moreover, we find situations where the Hamiltonian non-Hermiticity is small (due to a term with a large norm) while the non-Hermiticity score of simple observables is large, and vice versa. Thus, different non-Hermiticity quantifiers complement each other, being able to reveal different notions of non-Hermiticity. Furthermore, as we demonstrate in our work, the scores may become instrumental in marking non-Hermitian phase transitions.
@@
The rest of this article is organized as follows. In Sec. II, we introduce the formalism that provides an operational interpretation of dynamical evolution under any non-Hermitian Hamiltonian, unveil the physical meaning of the right and left eigenvectors associated with a non-Hermitian Hamiltonian, and introduce the quantifiers “Hamiltonian non-Hermiticity” as well as the “non-Hermiticity score.” In Sec. III, we illustrate the framework, first at the example of a non-Hermitian Hamiltonian constructed using a set of imperfect Bell states. Subsequently, we discuss the behavior in the interacting Hatano–Nelson model with asymmetric hopping, for which we compare non-Hermiticity of the model Hamiltonian and the non-Hermiticity score for single- as well as multi-site properties. Finally, in Sec. IV, we conclude with a brief discussion.
@@ II Theoretical background
In this section, we provide the theoretical framework for quantifying non-Hermiticity of a quantum system. We first discuss an operational interpretation of non-Hermitian dynamics as application of positive operator valued measurement on the system which is followed by a postselection. This interpretation shows that the resulting time-evolved state has a valid (albeit non-normalized) probability distribution, meaning observables can be measured and a physical meaning can be associated to them. Then, we show that the natural ensemble to evaluate expectation values is the ensemble of right and left eigenvectors for dynamics generated under $H_{\mathrm{nh}}$ and $H_{\mathrm{nh}}^{\dagger}$, respectively. Finally, this naturally leads us to two (not necessarily equivalent) ways to quantify non-Hermiticity of a quantum system: through the difference between the matrices $H_{\mathrm{nh}}$ and $H_{\mathrm{nh}}^{\dagger}$, and—perhaps more physically accessible—through the differences between expectation values reached by the evolution generated by $H_{\mathrm{nh}}$ and its hermitian conjugate.
II.1 Operational interpretation of dynamics under any non-Hermitian evolution
We start our discussion by providing an operational interpretation of dynamics governed by any generic non-Hermitian Hamiltonian $H_{\mathrm{nh}}$. Let us consider a general quantum state
$\rho(t)=\sum_{\lambda}p_{\lambda}\hat{P}_{\lambda},$ (1)
$\rho(t+\delta t)=\sum_{m}\mathcal{K}_{m}(\delta t)\rho(t)\mathcal{K}^{\dagger}_{m}(\delta t)=\sum_{\lambda}p_{\lambda}\hat{\tilde{P}}_{\lambda},$ (2)
where $\sum_{m}\mathcal{K}^{\dagger}_{m}\mathcal{K}_{m}=\mathbb{I}$ and $\hat{\tilde{P}}_{\lambda}=\sum_{m}\mathcal{K}_{m}(\delta t)L^{\dagger}_{\lambda}L_{\lambda}\mathcal{K}^{\dagger}_{m}(\delta t)$ is again a positive operator (because each summand $\mathcal{K}_{m}L^{\dagger}_{\lambda}L_{\lambda}\mathcal{K}^{\dagger}_{m}$ is), with $\hat{\tilde{P}}_{\lambda}\geq 0$ and $\sum_{\lambda}\hat{\tilde{P}}_{\lambda}=\mathbb{I}$.
We now consider a constrained evolution by projecting the evolved state $\rho(t+\delta t)$ onto a subspace,
$\tilde{\rho}(t+\delta t)=\mathcal{P}\rho(t+\delta t)\mathcal{P}=\sum_{\lambda}p_{\lambda}\mathcal{M}_{\lambda}\,,$ (3)
where $\mathcal{M}_{\lambda}=\mathcal{P}\hat{\tilde{P}}_{\lambda}\mathcal{P}$. In the case where $\mathcal{P}\neq\mathbb{I}$ (i.e., a projector onto a true subspace of the full Hilbert space), we have $\sum_{\lambda}\mathcal{M}_{\lambda}\neq\mathbb{I}$. The subspace restriction thus can lead to a loss of norm, $\mathrm{Tr}(\tilde{\rho}(t+\delta t))\leq 1$. Nevertheless, $\tilde{\rho}(t+\delta t)$ satisfies all the necessary conditions for a valid quantum state in the language of kinematic evolution:
  • (i) Positive operator: We have $\mathcal{M}_{\lambda}\geq 0$ and $p_{\lambda}\geq 0$. This implies $\tilde{\rho}(t+\delta t)\geq 0$.
  • (ii) Positive measurement probabilities: If we measure some quantum observable $\hat{O}=\sum_{\ell}a_{\ell}O_{\ell}$, the probabilities of distinct outcomes $a_{\ell}$ are positive, since using (i) we have $p_{\ell}\equiv\mathrm{Tr}(O_{\ell}\tilde{\rho}(t+\delta t)O_{\ell})\geq 0$.
Thus, while $\sum_{\ell}p_{\ell}=\mathrm{Tr}(\tilde{\rho}(t+\delta t))\leq 1$ is possible, probabilities of obtaining measurement results are always positive as expected from a physical observable. The remaining probability is associated to the cases where the system has leaked out into the subspace not considered (given by projector $\mathbb{I}-\mathcal{P}$). Hence, one can always assign an “operational meaning” or dynamical interpretation to any quantum state that satisfies the above two conditions. Before computing a quantity related to the system, it can be useful to normalize the quantum state $\tilde{\rho}(t+\delta t)$, as we do throughout this paper, so as to obtain the probabilities conditioned on the system being in the subspace, which sum up to identity.
We now show that the evolution under an arbitrary non-Hermitian Hamiltonian $H_{\mathrm{nh}}$, defined on the system’s Hilbert space $\mathcal{H}_{\mathrm{S}}$, can be written as a Kraus map acting on an extended space including an ancilla, $\mathcal{H}_{\mathrm{S}}\otimes\mathcal{H}_{\mathrm{A}}$, followed by a suitable projection. I.e., the evolution under any non-Hermitian Hamiltonian takes the form given by Eq. (3). The ancilla can be considered as any generic system ranging from qubits to bosonic baths. For the illustration purpose, here we consider it to be a qubit.
We start by constructing the evolution to leading order in a small time step $\delta t$, from which we obtain the continuous time evolution in the limit $\delta t\to 0$. Given any $H_{\mathrm{nh}}$, one can always define the operators $\mathcal{K}_{0}=(\mathbb{I}_{\mathrm{S}}-iH_{\mathrm{nh}}\delta t)\otimes\mathbb{I}_{\mathrm{A}}$ and $\mathcal{K}_{1}=\sqrt{\delta t(H_{\mathrm{nh}}^{\dagger}-H_{\mathrm{nh}}})\otimes\sigma_{\mathrm{A}}^{x}$ acting in $\mathcal{H}_{\mathrm{S}}\otimes\mathcal{H}_{\mathrm{A}}$. Here, $\sigma_{\mathrm{A}}^{x}$ is the Pauli matrix and $\mathcal{K}_{0}^{\dagger}\mathcal{K}_{0}+\mathcal{K}_{1}^{\dagger}\mathcal{K}_{1}=\mathbb{I}+\mathcal{O}(\delta t^{2})$. The dynamical evolution of an initial state operator $\rho(t)$ acting on $\mathcal{H}_{\mathrm{S}}\otimes\mathcal{H}_{\mathrm{A}}$ is given by
$\rho(t+\delta t)=\mathcal{K}_{0}\rho(t)\mathcal{K}_{0}^{\dagger}+\mathcal{K}_{1}\rho(t)\mathcal{K}_{1}^{\dagger}\,.$ (4)
This evolution describes a Master equation with jump operators $\mathcal{K}_{1}$ and is in the Kraus form as given in Eq. (2). (Note that the considered form of the operators $\mathcal{K}_{0}$ and $\mathcal{K}_{1}$ is not unique; in particular, a specific physical implementation yielding the same $H_{\mathrm{nh}}$ may be described microscopically by a different set of Kraus or unitary operators [24, 25, 26, 27, 28, 29].)
To obtain the desired evolution, we assume the ancilla qubit to be initialized in the $|0\rangle\langle 0|$ state. We then employ dynamical evolution under the above map. Afterwards, we project the evolved state back onto $\mathcal{P}_{0}=|0\rangle\langle 0|$. This amounts to using the qubit as a “flag” permitting postselection onto the no-jump trajectory [5, 30]. The result is
$\mathcal{P}_{0}\rho(t+\delta t)\mathcal{P}_{0}$ $=$ $\mathcal{P}_{0}\mathcal{K}_{0}(\rho_{\mathrm{S}}(t)\otimes|0\rangle\langle 0|)\mathcal{K}_{0}^{\dagger}\mathcal{P}_{0}$ (5)
$+\mathcal{P}_{0}\mathcal{K}_{1}(\rho_{\mathrm{S}}(t)\otimes|0\rangle\langle 0|)\mathcal{K}_{1}^{\dagger}\mathcal{P}_{0}$
$=$ $(\mathbb{I}_{\mathrm{S}}-iH_{\mathrm{nh}}\delta t)\rho_{\mathrm{S}}(t)(\mathbb{I}_{\mathrm{S}}+iH_{\mathrm{nh}}^{\dagger}\delta t)\otimes|0\rangle\langle 0|$
$+\mathcal{O}(\delta t^{2}).$
Therefore, in the limit of a small time step $\delta t$, the system undergoes the non-Hermitian evolution given by $e^{-itH_{\mathrm{nh}}}\rho_{\mathrm{S}}e^{itH^{\dagger}_{\mathrm{nh}}}$, whereas the ancilla remains in the initialized $|0\rangle\langle 0|$ state. One can see this evolution as a quantum-Zeno effect, generated by the continuous monitoring described by $\mathcal{P}_{0}$, and which freezes the ancilla to the $|0\rangle\langle 0|$ state. As is typical for the quantum-Zeno effect, the system undergoes a non-Hermitian evolution, describing the loss of norm out of the Zeno subspace [31, 32, 33].
In summary, if we consider a quantum system evolved under the effect of an arbitrary non-Hermitian system, we can always assign an operational meaning to the evolved states: they can be understood as quantum states originated due to the action of a constrained dynamical evolution manifested using complete quantum measurement and postselection of the results. This interpretation shows that the resulting time-evolved state has a valid probability distribution, meaning observables can be measured and a physical meaning can be associated to them.
II.2 Interpreting the eigenbasis of $H_{\mathrm{nh}}$
$H_{\mathrm{nh}}=\sum_{m}\lambda_{m}|R_{m}\rangle\langle L_{m}|,$ (6)
$\rho_{RR}(t)$ $=$ $e^{-iH_{\mathrm{nh}}t}\rho_{\mathrm{in}}e^{itH_{\mathrm{nh}}^{\dagger}},$ (7)
where $a_{mn}=\langle L_{m}|\rho_{\mathrm{in}}|L_{n}\rangle$. We can distinguish three situations.
Under this condition, Eq. (7) can be rewritten as
$\rho_{RR}(t)$ $=\sum_{m}e^{2t\Im[\lambda_{m}]}a_{mm}|R_{m}\rangle\langle R_{m}|$ (8)
At large times, $\rho_{RR}(t)$ eventually converges to the right eigenvector with slowest decay rate [5, 34],
$\rho_{RR}(t)\xrightarrow{t\rightarrow\infty}e^{2t\lambda}a_{\tilde{m}\tilde{m}}|R_{\tilde{m}}\rangle\langle R_{\tilde{m}}|.$ (9)
In general, physical expectation values are given by $\mathrm{Tr}(\rho_{RR}(t)\hat{O})$. Taking the trace in the biorthogonal basis that satisfies the completeness relation $\sum_{m}|R_{m}\rangle\langle L_{m}|=\mathbb{I}$, and using the relation $\langle R_{n}|L_{k}\rangle=\delta_{nk}$, we thus find physical observables will be obtained from the expectation values $\langle R_{\tilde{m}}|\hat{O}|R_{\tilde{m}}\rangle$.
2. All $\lambda_{m}$’s are real (Parity-Time symmetric case).
In this case, Eq. (7) reduces to
$\rho_{RR}(t)$ $=\sum_{m}a_{mm}|R_{m}\rangle\langle R_{m}|$
$\rho_{RR}(t)\xrightarrow{t\rightarrow\infty}\sum_{m}a_{mm}|R_{m}\rangle\langle R_{m}|.$ (11)
3. Arbitrary non-Hermitian time evolution.
In the general scenario, an observable $\hat{O}$ undergoes the evolution
$\text{Tr}(\hat{O}\rho_{RR}(t))$ $=$ $\sum_{mn}e^{-it(\lambda_{m}-\lambda_{n}^{*})}a_{mn}\langle R_{n}|\hat{O}|R_{m}\rangle\,.$ (12)
All the scenarios considered above illustrate the physical origin of expectation values evaluated in the right eigenvectors of the model. A similar analysis with $H_{\mathrm{nh}}^{\dagger}$ justifies the same reasoning for the left eigenvectors ($\{\langle L_{m}|\}$). As these considerations also show, while possibly containing interesting properties of the model [34], terms like $\langle R_{m}|\hat{O}|L_{n}\rangle$ do not naturally appear in the evaluation of physical quantities for the quantum state evolved under any generic non-Hermitian dynamics.
II.3 Non-Hermiticity quantifiers
@@
The above considerations put us into the position to introduce two (nonequivalent) ways of quantifying non-Hermiticity through the full Hamiltonian or through observables.
II.3.1 Hamiltonian non-Hermiticity

### v2

@@ Abstract
The non-Hermitian paradigm of quantum systems displays salient features drastically different from Hermitian counterparts. In this work, we focus on one such aspect, the difference of evolving quantum ensembles under $H_{\mathrm{nh}}$ (right ensemble) versus its Hermitian conjugate, $H_{\mathrm{nh}}^{\dagger}$ (left ensemble). We propose a formalism that quantifies the (dis-)similarity of these right and left ensembles, for single- as well as many-particle quantum properties. Such a comparison gives us a scope to measure the extent to which non-Hermiticity gets translated from the Hamiltonian into physically observable properties. We test the formalism in two cases: First, we construct a non-Hermitian Hamiltonian using a set of imperfect Bell states, showing that the non-Hermiticity of the Hamiltonian does not automatically comply with the non-Hermiticity at the level of observables. Second, we study the interacting Hatano–Nelson model with asymmetric hopping as a paradigmatic quantum many-body Hamiltonian. Interestingly, we identify situations where the measures of non-Hermiticity computed for the Hamiltonian, for single-, and for many-particle quantum properties behave distinctly from each other. Thus, different notions of non-Hermiticity can become useful in different physical scenarios. Furthermore, we demonstrate that the measures can qualitatively mark the model’s Parity–Time (PT) symmetry-breaking transition. Our findings can be instrumental in unveiling new exotic quantum phases of non-Hermitian quantum many-body systems as well as in preparing resourceful states for quantum technologies.
[…]
@@
Here, we present such a framework. To provide a concise overview, we briefly anticipate its main features, while we discuss its details in the bulk of the article. As is well known, the left and right eigenvectors diagonalizing a non-Hermitian Hamiltonian $\hat{H}_{\mathrm{nh}}=\sum_{m}\lambda_{m}|R_{m}\rangle\langle L_{m}|$, obeying $\hat{H}_{\mathrm{nh}}|R_{m}\rangle=\lambda_{m}|R_{m}\rangle$, $\langle L_{m}|\hat{H}_{\mathrm{nh}}=\langle L_{m}|\lambda_{m}$, and $\langle L_{m}|R_{n}\rangle=\delta_{mn}$, can in general be distinct, $\langle L_{m}|\neq(|R_{m}\rangle)^{\dagger}$. As we discuss, the evolution generated by $\hat{H}_{\mathrm{nh}}$ leads an initial state to ensembles of right eigenvectors (including, e.g., single eigenstates reachable at large times), $\rho_{RR}(t)=\frac{e^{-it\hat{H}_{\mathrm{nh}}}\rho_{\mathrm{in}}e^{it\hat{H}^{\dagger}_{\mathrm{nh}}}}{\mathrm{Tr}[e^{-it\hat{H}_{\mathrm{nh}}}\rho_{\mathrm{in}}e^{it\hat{H}^{\dagger}_{\mathrm{nh}}}]}$, such that observables measured on the system will physically be of the form $\langle R|\hat{O}|R\rangle$. In contrast, the evolution governed by $\hat{H}_{\mathrm{nh}}^{\dagger}$ leads to ensembles of left eigenvectors, $\rho_{LL}(t)=\frac{e^{-it\hat{H}^{\dagger}_{\mathrm{nh}}}\rho_{\mathrm{in}}e^{it\hat{H}_{\mathrm{nh}}}}{\mathrm{Tr}[e^{-it\hat{H}^{\dagger}_{\mathrm{nh}}}\rho_{\mathrm{in}}e^{it\hat{H}_{\mathrm{nh}}}]}$, and thus to expectation values of the form $\langle L|\hat{O}|L\rangle$. As we further show, the dynamics under any non-Hermitian Hamiltonian $\hat{H}_{\mathrm{nh}}$ can be modeled as an effect of complete quantum measurements and subsequent post-selection of the results, giving a clear physical meaning to these expectation values in terms of observable properties. Unlike the Hermitian case, the evolution governed by $\hat{H}_{\mathrm{nh}}$ and $\hat{H}_{\mathrm{nh}}^{\dagger}$ may lead to results that can significantly differ from each other, due to a rich interplay between potentially complex $\lambda_{m}$ and/or the difference between $\{|R_{m}\rangle\}$ and $\{\langle L_{m}|\}$. Our aim is to design measures that systematically capture these differences between $\hat{H}_{\mathrm{nh}}$ and $\hat{H}_{\mathrm{nh}}^{\dagger}$, and thus quantify the non-Hermiticity of the system through observable properties.
[…]
@@
Besides presenting complementary possibilities to certify non-Hermiticity, the above formalism opens up the possibility to analyze to what extent the degree of non-Hermiticity of the Hamiltonian $\hat{H}_{\mathrm{nh}}$ gets translated to the behavior of physically observable quantities. Moreover, we find situations where the Hamiltonian non-Hermiticity is small (due to a term with a large norm) while the non-Hermiticity score of simple observables is large, and vice versa. Thus, different non-Hermiticity quantifiers complement each other, being able to reveal different notions of non-Hermiticity. Furthermore, as we demonstrate in our work, the scores may become instrumental in marking non-Hermitian phase transitions.
@@
The rest of this article is organized as follows. In Sec. II, introduce the quantifiers “Hamiltonian non-Hermiticity” as well as the “non-Hermiticity score.” In Sec. III, we illustrate the framework, first at the example of a non-Hermitian Hamiltonian constructed using a set of imperfect Bell states. Subsequently, we discuss the behavior in the interacting Hatano–Nelson model with asymmetric hopping, for which we compare non-Hermiticity of the model Hamiltonian and the non-Hermiticity score for single- as well as multi-site properties. Finally, in Sec. IV, we conclude with a brief discussion.
@@ II Theoretical background
In this section, we introduce two (nonequivalent) ways of quantifying non-Hermiticity through the full Hamiltonian or through observables.
II.0.1 Hamiltonian non-Hermiticity